# Towards quantum-centric simulations of extended molecules: sample-based quantum diagonalization enhanced with density matrix embedding theory

### v1

@@ V Additional data
Figure 6: Sparsity of SQD wavefunctions for a hydrogen ring Hilbert space dimension (dashed line), number $d$ of configurations in the SQD wavefunction (symbols connected by a solid line) and number $d^{\prime}$ of such configurations with coefficients above $1\cdot 10^{-5}$ in absolute value squared (symbols connected by a dashed line), for a ring of 18 hydrogen atoms as a function of bondlength, studied with SQD (top, red squares) and DMET-SQD (bottom, green triangles).
@@
Figure 7: Sparsity of SQD wavefunctions for cyclohexane Hilbert space dimension (dashed line), number $d$ of configurations in the SQD wavefunction (triangles connected by a solid line) and number $d^{\prime}$ of such configurations with coefficients above $1.0\cdot 10^{-5}$ in absolute value squared (triangles connected by a dashed line), for cyclohexane conformations studied with DMET-SQD.

### v2

@@ V Additional data
Figure 6: Sparsity of SQD wavefunctions for a hydrogen ring Hilbert space dimension (dashed line), number $d$ of configurations in the SQD wavefunction (symbols connected by a solid line) and number $d^{\prime}$ of such configurations with coefficients above $1\cdot 10^{-8}$ in absolute value squared (symbols connected by a dashed line), for a ring of 18 hydrogen atoms as a function of bondlength, studied with SQD (top, red squares) and DMET-SQD (bottom, green triangles).
@@
Figure 7: Sparsity of SQD wavefunctions for cyclohexane Hilbert space dimension (dashed line), number $d$ of configurations in the SQD wavefunction (triangles connected by a solid line) and number $d^{\prime}$ of such configurations with coefficients above $1.0\cdot 10^{-8}$ in absolute value squared (triangles connected by a dashed line), for cyclohexane conformations studied with DMET-SQD.